# Gas meets Kozai: the influence of a gas-rich accretion disc on hierarchical triples undergoing von Zeipel-Lidov-Kozai oscillations

### v1

@@ Abstract
Active galactic nuclei (AGNs) consist of a central supermassive black hole (SMBH) embedded in a region with both high gas and stellar densities: the gas is present as a thin accretion disc that fuels the central SMBH, while the stars form a dense, roughly isotropic nuclear star cluster. The binaries present in such a cluster could be considered naturally as triples, with the SMBH as a third object, and their dynamics also depend on the interaction with the gas-rich disc. In this paper, we study the evolution of such a binary on an inclined orbit with respect to the disc. The binary experiences both eccentricity excitation via the von Zeipel-Lidov-Kozai (ZLK) effect and drag forces from each time it penetrates the disc. We find that, as the outer orbital inclination decreases, the evolution of inner orbital separation can transition from a regime of gradual hardening to a regime of rapid softening. As such binaries grow wider, their minimum pericentre distances (during ZLK oscillations) decrease. We show that a simple geometric condition, modulated by the complex ZLK evolution, dictates whether a binary expands or contracts due to the interactions with the AGN disc. Our results suggest that the interaction with gas-rich accretion disc could enhance the rate of stellar mergers and formation of gravitational wave sources, as well as other transients. The treatment introduced here is general and could apply, with the proper modifications, to hierarchical triples in other gas-rich systems.
@@
keyword1 – keyword2 – keyword3
@@ 1 Introduction
However, the standard ZLK treatment neglects other environmental effects, such as the presence of gas. Gas-rich systems are abundant on various scales, including protoplanetary discs, AGN discs, gas-rich star clusters, and the interstellar medium. The presence of gas can alter the dynamics of stellar, planetary and compact objects, leading to the formation of unique dynamics and transients (e.g., Bahcall & Ostriker, 1975; Artymowicz et al., 1993; Grishin & Perets, 2015, 2016; Fan et al., 2024; Kaur & Stone, 2025a; Rozner et al., 2025) that depend on the geometry and density of the gas distribution, including binary formation (e.g, Li et al., 2023; Rowan et al., 2023, 2024b; DeLaurentiis et al., 2023; Rozner et al., 2023; Whitehead et al., 2024a, b; Whitehead et al., 2025; Dodici & Tremaine, 2024), mergers (e.g., McKernan et al., 2012; Stone et al., 2017; Yang et al., 2019; Tagawa et al., 2020; Rozner & Perets, 2022; Xue et al., 2025) and affecting the binary population (e.g., McKernan et al., 2024; Rowan et al., 2024a; Rozner & Perets, 2024). The gravitational effect of a disc of gas on the ZLK effect has also been explored, by considering its modification of the system’s secular precession frequencies (e.g., Chen et al., 2013; Fan et al., 2024).
@@ 2 Equations of Motion
We consider the dynamics of a stellar binary in orbit about a central SMBH in the presence of an AGN disc (see an illustration in Fig. 1). The inner binary has component masses $m_{1}$ and $m_{2}$, radii $R_{1}$ and $R_{2}$, semi-major axis $a$, eccentricity $e$, inclination $i$ with respect to the disc normal, and true anomaly $f$. We generally track its secular orbital evolution by its orbital angular momentum vector $\boldsymbol{\mathbf{L}}=L\hat{\boldsymbol{\mathbf{l}}}$ and its eccentricity vector $\boldsymbol{\mathbf{e}}=e\hat{\boldsymbol{\mathbf{e}}}$. The two components of the outer binary are the inner binary and the SMBH with mass $m_{3}$. Its orbital elements are denoted $a_{\rm out}$, $e_{\rm out}$, $i_{\rm out}$ (also with respect to the disc normal), and its orbit is described by $\boldsymbol{\mathbf{L}}_{\rm out}=L_{\rm out}\hat{\boldsymbol{\mathbf{l}}}_{\rm out}$ and $\boldsymbol{\mathbf{e}}_{\rm out}=e_{\rm out}\hat{\boldsymbol{\mathbf{e}}}_{\rm out}$ with analogous meanings to the inner orbit’s properties. The AGN disc is oriented with its disc normal along $\hat{z}$ and aspect ratio $(H/r)\approx 1/200$. We assume that the binary initially satisfies $i_{\rm out}\gg(H/r)$, and we aim to model the effect of repeated passages through the AGN disc on the evolution of the hierarchical triple. Specifically, we consider three effects. The first is the secular evolution of the hierarchical triple, primarily due to the von Zeipel-Lidov-Kozai (ZLK) effect. The second is the effect of gas drag on both the inner and outer orbits. The third is the rapid precession of the outer orbit driven by the gravitational potential of the disc. We describe each of these in turn below.
@@ 2.1 The von Zeipel-Lidov-Kozai (ZLK) Effect
where $j_{\min}\equiv\sqrt{1-e_{\max}^{2}}$ and $e_{\max}$ is the maximum eccentricity attained by the inner orbit during a ZLK cycle. $e_{\max}$ depends on both the general-relativistic and rotationally driven apsidal precession (the precession driven by the tidal bulge is subdominant here) via the two dimensionless parameters (Liu et al., 2015):
ϵGR=subscriptitalic-$\epsilon_{\rm GR}={}$ $\frac{3Gm_{12}^{2}\tilde{a}_{\rm out}^{3}}{a_{\rm in}^{4}c^{2}m_{3}}$
[…]
$\epsilon_{\mathrm{rot},i}={}$ $\frac{m_{12}\tilde{a}_{\rm out}^{3}k_{2,i}R_{i}^{5}}{2Ga^{5}m_{k}m_{3}}\omega_{i}^{2},$ (5)
[…]
×$\times\left(\frac{R}{R_{\odot}}\right)^{2}\left(\frac{a_{\rm in}}{1\;\mathrm{au}}\right)^{-5}\left(\frac{m_{3}}{10^{7}M_{\odot}}\right)^{-1}\left(\frac{f_{i}}{0.3}\right)^{2},$ (6)
$\frac{9e_{\max}^{2}}{8}={}$ $\frac{\epsilon_{\rm rot,1}+\epsilon_{\rm rot,2}}{3}\left(\frac{1}{j_{\min}^{3}}-1\right)+\epsilon_{\rm GR}\left(\frac{1}{j_{\min}}-1\right),$ (7)
where $k\neq i$ indexes the companion star. ZLK is truncated when either $\epsilon_{\rm GR}>9/4$ or $\epsilon_{\rm rot,1}+\epsilon_{\rm rot,2}>9/4$ (Liu et al., 2015; Su, 2025). For the fiducial parameters, we find that ZLK suppression is dominated by the rotational SRF. This lets us rewrite the condition for double-averaging being valid as
@@
[…]
×$\times\left(\frac{R}{R_{\odot}}\right)^{2/3}\left(\frac{a_{\rm in}}{1\;\mathrm{au}}\right)^{-5/3}\left(\frac{m_{3}}{10^{7}M_{\odot}}\right)^{-1/3}\left(\frac{f_{i}}{0.3}\right)^{2/3},$ (8)
@@ 2.2 Effect of Gas on Binary Orbit
There are several models for the rate that an object decelerates when moving through a gaseous medium, including gas dynamical friction (Ostriker, 1999) and disc-like migration/formation of mini-discs (e.g., Artymowicz et al., 1991; McKernan et al., 2012; Stone et al., 2017; Tagawa et al., 2020). In this work, we adopt the so-called “aerodynamic” drag prescription (e.g., Artymowicz et al., 1993; Šubr & Karas, 1999; Generozov & Perets, 2022; Wang et al., 2024) given by
@@
where $R$ is the star’s physical radius and $c_{\rm s}$ is the local sound speed (e.g., a continuous version of the prescription adopted in Wang et al., 2024). For a sense of scale, we comment that
@@
$t_{\rm cross}$ $\simeq 185\;\mathrm{day}\frac{1}{\sin i_{\rm out}}\left(\frac{a_{\rm out}}{10^{4}\;\mathrm{au}}\right)^{1/2}\left(\frac{H/r}{0.005}\right),$ (14)
@@ 2.3 Disc-Driven Precession
Finally, due to the large mass of the disc, it will drive orbital precession of the outer orbit; the effect on the inner orbit is weaker by a factor of $(a/a_{\rm out})^{2}$ and is negligible. While this has been (reasonably) omitted in previous studies of interactions of single stars with AGN discs, it cannot be neglected here, since the outer orbit orientation plays a key role in the ZLK dynamics of the inner binary. We cannot directly use the expressions available in the literature (e.g., Terquem & Ajmia, 2010; Zhao et al., 2012; Chen et al., 2013) since they are derived in the limit where $a_{\rm out}\ll r_{\rm in}$, i.e. where the disc is exterior to the precessing orbit. In the case where the orbit and disc have comparable radii, the precession rate is formally divergent: multipolar expansions break down near $a_{\rm out}\simeq r$, and the Laplace coefficients in standard Laplace-Lagrange theory are also singular. Instead, a simple calculation (Appendix A.3) shows that the precession induced by a disc on an orbit with semi-major axis $a_{\rm out}$ depends on the gas density $\rho_{g}$ of the disc at $a_{\rm out}$ via:
$\frac{d\boldsymbol{\mathbf{L}}_{\rm out}}{dt}\simeq{}$ $\Omega_{\rm out}(\boldsymbol{\mathbf{L}}_{\rm out}\times\hat{\boldsymbol{\mathbf{z}}}),$ (16)
$\Omega_{\rm out}\sim$ $\sqrt{\pi G\rho_{g}(H/a)\cot i_{\rm out}\cos i_{\rm out}}\frac{\pi}{2}$
$\approx$ $\frac{2\pi}{2\;\mathrm{kyr}}\sqrt{\cot i_{\rm out}\cos i_{\rm out}}\left(\frac{\rho_{g}}{10^{11}\;M_{\odot}/\mathrm{pc}^{3}}\right)^{1/2}\left(\frac{H/r}{0.005}\right)^{1/2},$ (17)
and identically for $\boldsymbol{\mathbf{e}}_{\rm out}$. Especially as the outer orbit begins to align with the disc, this is multiple orders of magnitude faster than the next fastest timescale in the system. We adopt a fixed value of $\Omega_{\rm out}=2\pi/3\;\mathrm{kyr}$ for computational efficiency, as further increase of $g$ will not affect the dynamics.
@@ 3.1 Characteristic timescales
[…]
@@
[…]
@@ 3.2 A more realistic disc model
The ratio of the gaseous to ZLK timescale serves as a proxy for the level of influence we can expect gas effects to alter the ZLK evolution of the inner binary. Unsurprisingly, this closely follows the $\Sigma$ distribution. Considering binaries of equal hardness, i.e $a_{\text{in}}/r_{\text{H}}$ is constant, where $r_{\text{H}}=a_{\text{out}}(m_{12}/(3m_{3}))^{1/3}$ is the Hill radius, then under the assumption that comparatively few binaries will exist within $R<10^{-4}\ \rm{pc}$ we would expect binaries undergoing ZLK oscillations to be increasingly affected by gaseous drag. We find this result is largely independent of the choice of $m_{3}$, with $t_{\text{gas}}/t_{\text{ZLK}}$ ranging from $10^{3}$ to $<10^{1}$ for $R=10^{-1}\ \rm{pc}$ to $10^{-3}\ \rm{pc}$ for typical AGN SMBH masses.
@@ 4 Results
We consider a binary with $m_{1}=2M_{\odot}$ and $m_{2}=M_{\odot}$ in orbit around a SMBH with $m_{3}=10^{7}M_{\odot}$. For ease of comparison, we take the separation of the inner orbit to be $a=1\;\mathrm{au}$ and of the outer orbit to be $a_{\rm out}=10^{4}\;\mathrm{au}$, but we allow the disc surface density to be a free parameter. We truncate the system evolution once $i_{\rm out}$ decreases below $(H/r)=1/200$, whereupon the binary becomes fully embedded in the disc. The results of such an evolution are shown for four gas surface densities in Fig. 3, where the disc surface density is reported at $a_{\rm out}$; we take $\Sigma=\Sigma(a_{\rm out})(a_{\rm out}/r_{\rm out})$, i.e. a $p=-1$ surface density profile. We see that the outer orbit evolves in the expected manner, where the orbit shrinks, circularizes, and aligns with the disc (e.g., Wang et al., 2024).
Figure 3: Evolution of an unequal mass-ratio system ($m_{1}=2M_{\odot}$ and $m_{2}=M_{\odot}$) with both stars rotating at $\omega_{i}=0.3\sqrt{Gm_{i}/R_{i}^{3}}$ for four different gas densities (the value in the legend corresponds to the density at $a_{\rm out}$, and we take $\Sigma(r)=\Sigma(a_{\rm out})a_{\rm out}/r$); we use $a=1\;\mathrm{au}$, $a_{\rm out}=10^{4}\;\mathrm{au}$, $e_{\rm out}=0.1$, $i_{\rm out}=30^{\circ}$, and we take initial values $e_{0}=0.2$ and $i_{0}=80^{\circ}$. The left column of panels shows the evolution of the inner orbit: the top panel illustrates that the orbit eventually softens, particularly at late times. The second and third panels show the signature ZLK eccentricity-inclination cycles, though the inclination is given with respect to $\hat{\boldsymbol{\mathbf{z}}}$, the disc normal. The ¿stochasticity in the eccentricity and inclination cycles is due to the slight aperiodicity introduced by the finite precession rate of the outer orbit driven by the disc. The fourth panel shows the instantaneous $\dot{a}$, where significant structure can be seen; see detailed discussion in Section 4.1. The final panel shows the inner orbit’s argument of pericentre $\omega$ with respect to the disc plane. The first three panels of the right column show the evolution of the outer orbit: the orbit shrinks, circularizes, and aligns with the disc plane. The fourth panel shows the ratio of the geometric and BHL contributions to the drag force cross-section (Eq. 12). The fifth panel shows the ratio of the orbit’s pericentre distance to the radii of the stars; recall that the Roche radius of a star is $\sim 2R_{\star}$.
@@
On the other hand, the inner orbit undergoes several phenomena worth discussing in detail. First, note that the inner orbit executes ZLK oscillations about $\hat{z}$ instead of $\hat{l}_{\rm out}$ (bottom-right panels of Fig. 3). This is because $\hat{l}_{\rm out}$ is rapidly varying, with average orientation along $\hat{z}$. Second, note that the inner orbit’s semi-major axis evolution broadly transitions from inspiral to outspiral as the outer orbital inclination decreases. While the outspiraling regime has been previously characterized (Sánchez-Salcedo & Chametla, 2014; Grishin & Perets, 2016), its dependence on the outer orbit’s geometry is novel to this work. For comparison, we also display the evolution of an equal mass ratio binary in Fig. 4, with all other parameters held constant; no significant qualitative changes are seen in the evolution.

### v2

@@ Abstract
Active galactic nuclei (AGNs) consist of a central supermassive black hole (SMBH) embedded in a region with both high gas and stellar densities: the gas is present as a thin accretion disc that fuels the central SMBH, while the stars form a dense, roughly isotropic nuclear star cluster. The binaries present in such a cluster could be considered naturally as triples, with the SMBH as a third object, and their dynamics also depend on the interaction with the gas-rich disc. In this paper, we study the evolution of such a binary on an inclined orbit with respect to the disc. The binary experiences both eccentricity excitation via the von Zeipel-Lidov-Kozai (ZLK) effect and drag forces from each time it penetrates the disc. We find that, as long as the inner binary remains in the ZLK regime, then the evolution of inner orbital separation can transition from a regime of gradual hardening to a regime of rapid softening as the outer orbital inclination decreases. As such binaries grow wider, their minimum pericentre distances (during ZLK oscillations) decrease. We show that a simple geometric condition, modulated by the complex ZLK evolution, dictates whether a binary expands or contracts due to the interactions with the AGN disc. Our results suggest that the interaction with gas-rich accretion disc could enhance the rate of stellar mergers and formation of gravitational wave sources, as well as other transients. The treatment introduced here is general and could apply, with the proper modifications, to hierarchical triples in other gas-rich systems.
@@
Accretion discs – stars:binaries – galaxies:nuclei
@@ 1 Introduction
However, the standard ZLK treatment neglects other environmental effects, such as the presence of gas. Gas-rich systems are abundant on various scales, including protoplanetary discs, AGN discs, gas-rich star clusters, and the interstellar medium. The presence of gas can alter the dynamics of stellar, planetary and compact objects, leading to the formation of unique dynamics and transients (e.g., Bahcall & Ostriker, 1975; Artymowicz et al., 1993; Grishin & Perets, 2015, 2016; Fan et al., 2024; Kaur & Stone, 2025a; Rozner et al., 2025) that depend on the geometry and density of the gas distribution, including binary formation (e.g, Li et al., 2023; Rowan et al., 2023, 2024b; DeLaurentiis et al., 2023; Rozner et al., 2023; Whitehead et al., 2024a, b, 2025; Dodici & Tremaine, 2024), mergers (e.g., McKernan et al., 2012; Stone et al., 2017; Yang et al., 2019; Tagawa et al., 2020; Rozner & Perets, 2022; Xue et al., 2025) and affecting the binary population (e.g., McKernan et al., 2024; Rowan et al., 2024a; Rozner & Perets, 2024). The gravitational effect of a disc of gas on the ZLK effect has also been explored, by considering its modification of the system’s secular precession frequencies (e.g., Chen et al., 2013; Fan et al., 2024).
@@ 2 Equations of Motion
We consider the dynamics of a stellar binary in orbit about a central SMBH in the presence of an AGN disc (see an illustration in Fig. 1). The inner binary has component masses $m_{1}$ and $m_{2}$, radii $R_{1}$ and $R_{2}$, semi-major axis $a$, eccentricity $e$, inclination $i$ with respect to the disc normal, and true anomaly $f$. We generally track its secular orbital evolution by its orbital angular momentum vector $\boldsymbol{\mathbf{L}}=L\hat{\boldsymbol{\mathbf{l}}}$ and its eccentricity vector $\boldsymbol{\mathbf{e}}=e\hat{\boldsymbol{\mathbf{e}}}$. The two components of the outer binary are the inner binary and the SMBH with mass $m_{3}$. Its orbital elements are denoted $a_{\rm out}$, $e_{\rm out}$, $i_{\rm out}$ (also with respect to the disc normal), and its orbit is described by $\boldsymbol{\mathbf{L}}_{\rm out}=L_{\rm out}\hat{\boldsymbol{\mathbf{l}}}_{\rm out}$ and $\boldsymbol{\mathbf{e}}_{\rm out}=e_{\rm out}\hat{\boldsymbol{\mathbf{e}}}_{\rm out}$ with analogous meanings to the inner orbit’s properties. The AGN disc is oriented with its disc normal along $\hat{z}$ and aspect ratio $h\approx 1/200$. We assume that the binary initially satisfies $i_{\rm out}\gg h$, and we aim to model the effect of repeated passages through the AGN disc on the evolution of the hierarchical triple. Specifically, we consider three effects. The first is the secular evolution of the hierarchical triple, primarily due to the von Zeipel-Lidov-Kozai (ZLK) effect. The second is the effect of gas drag on both the inner and outer orbits. The third is the rapid precession of the outer orbit driven by the gravitational potential of the disc. We describe each of these in turn below.
@@ 2.1 The von Zeipel-Lidov-Kozai (ZLK) Effect
where $j_{\min}\equiv\sqrt{1-e_{\max}^{2}}$ and $e_{\max}$ is the maximum eccentricity attained by the inner orbit during a ZLK cycle. $e_{\max}$ depends on both the general-relativistic and rotationally driven apsidal precession via the two dimensionless parameters (Liu et al., 2015) [^1] For distant binaries that are not tidally locked, the tidal bulge is much smaller than the rotational bulge for rapidly-rotating bodies (e.g. Liu et al., 2015). Note that while $\epsilon_{\rm rot,i}$ is developed for orbit-aligned rotational bulges, the effect of spin-orbit misalignment contributes only an order-unity correction (Correia et al., 2011).:
$\epsilon_{\rm GR}={}$ $\frac{3Gm_{12}^{2}\tilde{a}_{\rm out}^{3}}{a_{\rm in}^{4}c^{2}m_{3}}$
$={}$ $0.03\left(\frac{m_{\rm 12}}{3M_{\odot}}\right)^{2}\left(\frac{\tilde{a}_{\rm out}}{10^{4}\;\mathrm{au}}\right)^{3}\left(\frac{a_{\rm in}}{1\mathrm{au}}\right)^{-4}\left(\frac{m_{3}}{10^{7}M_{\odot}}\right)^{-1}.$ (4)
$\epsilon_{\mathrm{rot},i}={}$ $\frac{m_{12}\tilde{a}_{\rm out}^{3}k_{2,i}R_{i}^{5}}{2Ga^{5}m_{i}m_{3}}\omega_{i}^{2},$ (5)
$={}$ $0.012\left(\frac{k_{2,i}}{0.04}\right)\left(\frac{m_{12}}{3M_{\odot}}\right)\left(\frac{\tilde{a}_{\rm out}}{10^{4}\;\mathrm{au}}\right)^{3}$
$\times\left(\frac{R}{R_{\odot}}\right)^{2}\left(\frac{a_{\rm in}}{1\;\mathrm{au}}\right)^{-5}\left(\frac{m_{3}}{10^{7}M_{\odot}}\right)^{-1}\left(\frac{f_{i}}{0.3}\right)^{2},$ (6)
$\frac{9e_{\max}^{2}}{8}={}$ $\frac{\epsilon_{\rm rot,1}+\epsilon_{\rm rot,2}}{3}\left(\frac{1}{j_{\min}^{3}}-1\right)+\epsilon_{\rm GR}\left(\frac{1}{j_{\min}}-1\right).$ (7)
ZLK is truncated when either $\epsilon_{\rm GR}>9/4$ or $\epsilon_{\rm rot,1}+\epsilon_{\rm rot,2}>9/4$ (Liu et al., 2015; Su, 2025). For the fiducial parameters, we find that ZLK suppression is dominated by the rotational short-range forces. This lets us rewrite the condition for double-averaging being valid as
@@
$\simeq{}$ $0.2t_{\rm ZLK}\left(\frac{k_{2,i}}{0.04}\right)^{1/3}\left(\frac{m_{12}}{3M_{\odot}}\right)^{1/3}\left(\frac{\tilde{a}_{\rm out}}{10^{4}\;\mathrm{au}}\right)$
$\times\left(\frac{R}{R_{\odot}}\right)^{2/3}\left(\frac{a_{\rm in}}{1\;\mathrm{au}}\right)^{-5/3}\left(\frac{m_{3}}{10^{7}M_{\odot}}\right)^{-1/3}\left(\frac{f_{i}}{0.3}\right)^{2/3},$ (8)
@@ 2.2 Effect of Gas on Binary Orbit
There are several models for the rate that an object decelerates when moving through a gaseous medium, including gas dynamical friction (e.g., Ostriker, 1999; O’Neill et al., 2024) and disc-like migration/formation of mini-discs (e.g., Artymowicz et al., 1991; McKernan et al., 2012; Stone et al., 2017; Tagawa et al., 2020). In this work, we adopt the so-called “aerodynamic” drag prescription (e.g., Artymowicz et al., 1993; Šubr & Karas, 1999; Generozov & Perets, 2022; Wang et al., 2024) given by
@@
where $R$ is the star’s physical radius and $c_{\rm s}$ is the local sound speed (e.g., a continuous version of the prescription adopted in Wang et al., 2024)333 Note that $A\ll\pi r_{\rm H}^{2}$ for binaries crossing the disk on such inclined trajectories, where $r_{\rm H}=a_{\text{out}}(m_{12}/(3m_{3}))^{1/3}$ is the Hill radius. As such, the scale of the accretion flow is set by the BHL cross section $A$, within which the SMBH gravity is negligible.. For a sense of scale, we comment that
@@
$t_{\rm cross}$ $\simeq 185\;\mathrm{day}\frac{1}{\sin i_{\rm out}}\left(\frac{a_{\rm out}}{10^{4}\;\mathrm{au}}\right)^{1/2}\left(\frac{h}{0.005}\right),$ (14)
@@ 2.3 Disc-Driven Precession
Finally, due to the large mass of the disc, it will drive orbital precession of the outer orbit; the effect on the inner orbit is weaker by a factor of $(a/a_{\rm out})^{2}$ and is negligible. While this has been (reasonably) omitted in previous studies of interactions of single stars with AGN discs, it cannot be neglected here, since the outer orbit orientation plays a key role in the ZLK dynamics of the inner binary. There are two classes of expressions for this precession effect available in the literature, but neither are appropriate here: some (e.g., Terquem & Ajmia, 2010; Zhao et al., 2012; Chen et al., 2013) are derived in the limit where $a_{\rm out}\ll r_{\rm in}$ (i.e., where the disc is exterior to the precessing orbit), while others (e.g. Ward, 1981; Hahn, 2003; Zanazzi & Lai, 2018) assume that the orbital inclination is smaller than the disk aspect ratio (i.e., that the orbit is fully embedded in the disk). Instead, we show in Appendix A.3 that the precession induced by a disc on an orbit with semi-major axis $a_{\rm out}$ depends on the gas density $\rho_{g}$ of the disc at $a_{\rm out}$ via:
[…]
$\Omega_{\rm out}\simeq{}$ $\frac{G\rho_{g}}{n_{\rm out}}\frac{4h}{\sin i_{\rm out}}$
$={}$ $\frac{2\pi}{10\;\mathrm{kyr}}\frac{1}{\sin i_{\rm out}}\left(\frac{\rho_{g}}{10^{11}\;\mathrm{M_{\odot}/pc^{3}}}\right)\left(\frac{h}{0.005}\right)$
$\times\left(\frac{m_{3}}{10^{7}M_{\odot}}\right)^{-1/2}\left(\frac{a_{\rm out}}{10^{4}\;\mathrm{au}}\right)^{3/2},$ (17)
and identically for $\boldsymbol{\mathbf{e}}_{\rm out}$, where $h\equiv H/r$ is the disk aspect ratio. Especially as the outer orbit begins to align with the disc, this is substantially faster than the next fastest timescale in the system. We adopt a fixed value of $\Omega_{\rm out}=2\pi/10\;\mathrm{kyr}$ for computational efficiency, as further increase of $\Omega_{\rm out}$ will not affect the dynamics.
@@ 3.1 Characteristic timescales
$\approx{}$ $0.017\left(\frac{\rho_{g}}{10^{11}M_{\odot}/\mathrm{pc}^{3}}\right)\left(\frac{a_{\rm out}}{10^{4}\;\mathrm{au}}\right)^{2.5}\left(\frac{a}{1\;\mathrm{au}}\right)^{-1.5},$ (18)
@@
$={}$ $20\;\mathrm{Myr}\left(\frac{\rho_{g}}{10^{11}M_{\odot}/\mathrm{pc}^{3}}\right)^{-1}\left(\frac{a_{\rm out}}{10^{4}\;\mathrm{au}}\right)^{1/2}\left(\frac{\mu_{\rm out}}{3M_{\odot}}\right).$ (19)
@@ 3.2 A more realistic disc model
The ratio of the gaseous to ZLK timescale serves as a proxy for the level of influence we can expect gas effects to alter the ZLK evolution of the inner binary. Unsurprisingly, this closely follows the $\Sigma$ distribution. Considering binaries of equal hardness, i.e $a_{\text{in}}/r_{\text{H}}$ is constant, where $r_{\text{H}}$ is again the Hill radius, then under the assumption that comparatively few binaries will exist within $R<10^{-4}\ \rm{pc}$ we would expect binaries undergoing ZLK oscillations to be increasingly affected by gaseous drag. We find this result is largely independent of the choice of $m_{3}$, with $t_{\text{gas}}/t_{\text{ZLK}}$ ranging from $10^{3}$ to $10^{1}$ for $R=10^{-1}\ \rm{pc}$ to $10^{-3}\ \rm{pc}$ for typical AGN SMBH masses.
@@ 4 Results
We consider a binary with $m_{1}=2M_{\odot}$ and $m_{2}=M_{\odot}$ in orbit around a SMBH with $m_{3}=10^{7}M_{\odot}$. For ease of comparison, we take the separation of the inner orbit to be $a=1\;\mathrm{au}$ and of the outer orbit to be $a_{\rm out}=10^{4}\;\mathrm{au}$, but we allow the disc surface density to be a free parameter. We truncate the system evolution once $i_{\rm out}$ decreases below $h=1/200$, whereupon the binary becomes fully embedded in the disc. The results of such an evolution are shown for four gas surface densities in Fig. 3, where the disc surface density is reported at $a_{\rm out}$; we take $\Sigma=\Sigma(a_{\rm out})(a_{\rm out}/r_{\rm out})$, i.e. a p=1p=-1 surface density profile. We see that the outer orbit evolves in the expected manner, where the orbit shrinks, circularizes, and aligns with the disc (e.g., Wang et al., 2024).
Figure 3: Evolution of an unequal mass-ratio system ($m_{1}=2M_{\odot}$ and $m_{2}=M_{\odot}$) with both stars rotating at $\omega_{i}=0.3\sqrt{Gm_{i}/R_{i}^{3}}$ for four different gas densities (the value in the legend corresponds to the density at $a_{\rm out}$, and we take $\Sigma(r)=\Sigma(a_{\rm out})a_{\rm out}/r$); we use $a=1\;\mathrm{au}$, $a_{\rm out}=10^{4}\;\mathrm{au}$, $e_{\rm out}=0.1$, $i_{\rm out}=30^{\circ}$, and we take initial values $e_{0}=0.2$ and $i_{0}=80^{\circ}$. The left column of panels shows the evolution of the inner orbit: the top panel illustrates that the orbit eventually softens, particularly at late times. The second and third panels show the signature ZLK eccentricity-inclination cycles, though the inclination is given with respect to $\hat{\boldsymbol{\mathbf{z}}}$, the disc normal. The stochasticity in the eccentricity and inclination cycles is due to the slight aperiodicity introduced by the finite precession rate of the outer orbit driven by the disc. The fourth panel shows the instantaneous $\dot{a}$, where significant structure can be seen; see detailed discussion in Section 4.1. The final panel shows the inner orbit’s argument of pericentre $\omega$ with respect to the disc plane. The first three panels of the right column show the evolution of the outer orbit: the orbit shrinks, circularizes, and aligns with the disc plane. The fourth panel shows the ratio of the geometric and BHL contributions to the drag force cross-section (Eq. 12). The fifth panel shows the ratio of the orbit’s pericentre distance to the radii of the stars; recall that the Roche radius of a star is $\sim 2R_{\star}$.
@@
On the other hand, the inner orbit undergoes several phenomena worth discussing in detail. First, note that the inner orbit executes ZLK oscillations about $\hat{z}$ instead of $\hat{l}_{\rm out}$ (bottom-right panels of Fig. 3). This is because $\hat{l}_{\rm out}$ is rapidly varying, with average orientation along $\hat{z}$. Second, note that the inner orbit’s semi-major axis evolution broadly transitions from inspiral to outspiral as the outer orbital inclination decreases. While the outspiraling regime has been previously characterized (Sánchez-Salcedo & Chametla, 2014; Grishin & Perets, 2016), its dependence on the outer orbit’s geometry (when in the ZLK regime) is novel to this work. For comparison, we also display the evolution of an equal mass ratio binary in Fig. 4, with all other parameters held constant; no significant qualitative changes are seen in the evolution.